OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2. y - 5 = 3/2(x + 1) y + 1 = 3/2(x - 5) y + 5 = 3/2(x - 1) y - 1 = 3/2(x + 5)

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (anonymous):

I'm guessing D or C

OpenStudy (anonymous):

use this eq.. (y-y1)=m(x-x1) m=slope..compare your choice by putting values int is eq you will get result..

jigglypuff314 (jigglypuff314):

(-5, 1) as (x1, y1) into y - y1 = m(x - x1) you're close :)

OpenStudy (anonymous):

ok but dont you mean y^1 and x^1

jigglypuff314 (jigglypuff314):

ah no sorry for the confusion I mean \(y_1~~ and~~x_1\) subscripts

jigglypuff314 (jigglypuff314):

it is to tell the difference between y and \(y_1\) y - \(y_1\) = m(x - \(x_1\)) plug in (-5, 1) as (\(x_1\), \(y_1\)) that better?

OpenStudy (anonymous):

im not thinking at all today, this is what im writing (-51, 11) -_-

OpenStudy (anonymous):

i know im wrong, lol

jigglypuff314 (jigglypuff314):

lol xD y - (1) = m ( x - (-5)) is what I meant :P

OpenStudy (anonymous):

oh wait i think i understand a bit better, is that it? or is there more to it?

jigglypuff314 (jigglypuff314):

well you're gonna have to choose the letter that matches it right? ;) and negative times a negative makes a positive

OpenStudy (anonymous):

oh ok, thanks!